# PiCO: Peer Review in LLMs based on the Consistency Optimization

### v2

@@ 3 Experiments
3.1 Experimental Settings
Datasets. To validate the effectiveness of the proposed approach, we perform experiments on Chatbot Arena[59], MT-Bench[59], and AlpacaEval[32]. The MT-Bench dataset assesses six LLMs’ responses to 80 multi-category questions. The Chatbot Arena Conversations Dataset, with 33K conversations from 13K IPs during April-June 2023, evaluates real dialogue performance. AlpacaEval dataset integrates 805 evaluations from diverse tests (e.g., Self-Instruct[51], OASST, Anthropic’s helpful[7], Vicuna[16] and Koala[27] test sets) to align evaluations real-world interactions[22]. These datasets are collected by crowdsourcing platforms from human feedback, so they have a ground-truth ranking LLMs $\mathcal{R}^{*}$ aligned with human preferences.
LLMs Candidate Pool. In our experiments, we employ 15 LLMs with diverse architectures to construct the LLMs candidate pool, including GPT-3.5-Turbo[39], WizardLM-13B[55], Guanaco-33B[1], Vicuna-7B[16], Vicuna-13B[16], Koala-13B[26], Mpt-7B[46], gpt4all-13B[6], ChatGLM-6B[57], Oasst-sft-4-pythia-12B[20], FastChat-T5-3B[59], StableLM-7B[3], Dolly-12B[18], LLaMA-13B[47], Alpaca-13B[45]. All models use the same evaluation template as follows, more details can be found in Appendix B.
Template for Peer Review Stage System prompt: Please act as a judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You do not need to explain, just give your judgment. Output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better, and "[[C]]" for a tie. User Question: {question} Assistant A’s Answer: {answer a} Assistant B’s Answer: {answer b}
  • The wisdom of the crowds: The two methods that perform LLMs evaluation based on the wisdom of the crowds [44; 13; 54] are compared in this experiment. 1) Majority Voting [44]: Multiple review models vote for the better answer for the same response pair, and the model with the most votes gets 1 score; 2) Rating Voting [5]: Multiple review models also vote on the same response pair, and the number of votes obtained is the score.
  • State-of-the-art methods: The four recent SOTA methods of using either single or multiple models for self-evaluation are compared in this experiment. PandaLM[52]: It is a fine-tuned language model based on Llama-7b designed for the preference judgment tasks to evaluate and optimize LLMs. GPTScore[25]: It employs generative pre-trained models to assess the quality of generated text. It calculates the likelihood that the text was generated in response to specific instructions and context, indicative of high quality. In our implementation, GPT-3 (davinci-002) and flan-t5-xxl serve as the base models. PRD[31]: It transforms the LLMs win rates into weights for competitive ranking, while evaluating each LLM based on its preference for all possible pairs of answers, enabling a tournament-style ranking system. PRE[17]: It employs a supervised process to evaluate LLMs using a qualification exam, aggregates their scores based on accuracy, and assigns weights accordingly. PiCO (Ours): the proposed approach in this paper.
Metrics. For all experiments, we employ three metrics to evaluate the aforementioned experimental setups and our Peer Review method: PEN (Permutation Entropy), CIN (Count Inversions) and LIS (Longest Increasing Subsequence). Moreover, we perform the experiments for 4 runs and record the average results over 4 seeds ($seed=1,2,3,4$).
Table 1: The overall performance of our proposed method and the baselines under data volumes of 1, 0.7 and 0.4 on Chatbot Arena, MT-Bench, and AlpacaEval. A lower PEN score indicates better performance, as does a lower CIN score, while a higher LIS score signifies improved performance.
@@ 3.1 Experimental Settings
PEN ()($\downarrow$)( ↓ )
Majority Voting [44] $1.27^{\pm 0.05}$ $1.30^{\pm 0.03}$ $1.36^{\pm 0.06}$ $1.37^{\pm 0.03}$ $1.30^{\pm 0.06}$ $1.27^{\pm 0.04}$ $1.26^{\pm 0.02}$ $1.28^{\pm 0.03}$ $1.29^{\pm 0.03}$
Rating Voting [5] $1.39^{\pm 0.02}$ $1.43^{\pm 0.03}$ $1.42^{\pm 0.07}$ $1.32^{\pm 0.03}$ $1.35^{\pm 0.04}$ $1.38^{\pm 0.04}$ $1.34^{\pm 0.03}$ $1.37^{\pm 0.03}$ $1.34^{\pm 0.08}$
GPTScore(flan-t5-xxl)[25] $1.68^{\pm 0.01}$ $1.68^{\pm 0.02}$ $1.65^{\pm 0.02}$ $1.72^{\pm 0.02}$ $1.70^{\pm 0.02}$ $1.68^{\pm 0.03}$ $1.55^{\pm 0.02}$ $1.57^{\pm 0.03}$ $1.60^{\pm 0.01}$
GPTScore(davinci-002)[25] $1.54^{\pm 0.02}$ $1.64^{\pm 0.02}$ $1.68^{\pm 0.05}$ $1.51^{\pm 0.02}$ $1.61^{\pm 0.01}$ $1.61^{\pm 0.04}$ $1.25^{\pm 0.02}$ $1.23^{\pm 0.08}$ $1.26^{\pm 0.14}$
PandaLM[52] $1.65^{\pm 0.01}$ $1.64^{\pm 0.02}$ $1.63^{\pm 0.05}$ $1.55^{\pm 0.03}$ $1.59^{\pm 0.05}$ $1.52^{\pm 0.08}$ $1.56^{\pm 0.01}$ $1.58^{\pm 0.01}$ $1.64^{\pm 0.05}$
PRD[31] $1.15^{\pm 0.04}$ $1.12^{\pm 0.05}$ $1.13^{\pm 0.06}$ $1.15^{\pm 0.05}$ $1.17^{\pm 0.06}$ $1.23^{\pm 0.04}$ $1.21^{\pm 0.04}$ $1.22^{\pm 0.06}$ $1.23^{\pm 0.07}$
PRE[17] $1.07^{\pm 0.01}$ $1.03^{\pm 0.03}$ $1.06^{\pm 0.04}$ $1.17^{\pm 0.04}$ $1.13^{\pm 0.05}$ $1.19^{\pm 0.05}$ $1.18^{\pm 0.03}$ $1.21^{\pm 0.04}$ $1.15^{\pm 0.05}$
@@
CIN ()($\downarrow$)( ↓ )
Majority Voting [44] $22.00^{\pm 0.00}$ $23.25^{\pm 1.09}$ $25.00^{\pm 2.55}$ $23.00^{\pm 0.00}$ $20.50^{\pm 0.87}$ $21.00^{\pm 1.00}$ $20.00^{\pm 0.00}$ $21.25^{\pm 1.30}$ $22.25^{\pm 1.30}$
Rating Voting [5] $24.00^{\pm 0.00}$ $24.50^{\pm 1.29}$ $25.00^{\pm 1.15}$ $22.00^{\pm 0.00}$ $22.50^{\pm 1.00}$ $24.25^{\pm 0.50}$ $22.00^{\pm 0.00}$ $22.50^{\pm 0.58}$ $22.50^{\pm 1.00}$
GPTScore(flan-t5-xxl)[25] $67.00^{\pm 0.00}$ $66.50^{\pm 0.50}$ $68.25^{\pm 1.09}$ $53.00^{\pm 0.00}$ $55.75^{\pm 2.77}$ $54.50^{\pm 2.29}$ $35.00^{\pm 0.00}$ $36.00^{\pm 0.71}$ $37.75^{\pm 1.60}$
GPTScore(davinci-002)[25] $42.00^{\pm 0.00}$ $45.50^{\pm 1.12}$ $51.00^{\pm 5.61}$ $33.00^{\pm 0.00}$ $35.00^{\pm 0.71}$ $36.25^{\pm 1.64}$ $21.00^{\pm 0.00}$ $20.25^{\pm 2.86}$ $21.50^{\pm 4.39}$
PandaLM[52] $37.00^{\pm 0.00}$ $36.25^{\pm 1.79}$ $36.00^{\pm 3.74}$ $32.00^{\pm 0.00}$ $33.00^{\pm 3.32}$ $31.50^{\pm 6.34}$ $31.00^{\pm 0.00}$ $32.25^{\pm 1.30}$ $35.50^{\pm 2.60}$
PRD[31] $17.00^{\pm 0.00}$ $16.25^{\pm 0.43}$ $17.50^{\pm 1.50}$ $17.00^{\pm 0.00}$ $17.75^{\pm 1.09}$ $19.50^{\pm 1.50}$ $19.00^{\pm 0.00}$ $19.25^{\pm 1.48}$ $19.50^{\pm 0.87}$
PRE[17] $15.00^{\pm 0.00}$ $14.25^{\pm 0.83}$ $14.75^{\pm 1.09}$ $17.00^{\pm 0.00}$ $17.00^{\pm 1.00}$ $18.25^{\pm 1.30}$ $19.00^{\pm 0.00}$ $19.25^{\pm 1.09}$ $17.75^{\pm 1.30}$
PiCO (Ours) $\bm{12.00^{\pm 0.00}}$ $\bm{12.50^{\pm 0.50}}$ $\bm{12.25^{\pm 1.09}}$ $\bm{14.50^{\pm 0.50}}$ $\bm{14.75^{\pm 1.64}}$ $\bm{16.00^{\pm 6.36}}$ $\bm{17.00^{\pm 0.00}}$ $\bm{18.00^{\pm 1.87}}$ $\bm{17.25^{\pm 1.09}}$
LIS ()($\uparrow$)( ↑ )
Majority Voting [44] $7.00^{\pm 0.00}$ $6.75^{\pm 0.43}$ $6.75^{\pm 0.43}$ $7.00^{\pm 0.00}$ $\underline{8.25}^{\pm 0.43}$ $\underline{8.50}^{\pm 1.12}$ $8.00^{\pm 0.00}$ $7.50^{\pm 0.50}$ $7.50^{\pm 0.50}$
Rating Voting [5] $7.00^{\pm 0.00}$ $7.50^{\pm 0.58}$ $7.75^{\pm 0.50}$ $7.00^{\pm 0.00}$ $7.25^{\pm 0.50}$ $7.25^{\pm 0.50}$ $8.00^{\pm 0.00}$ $8.00^{\pm 0.00}$ $8.00^{\pm 0.00}$
GPTScore(flan-t5-xxl)[25] $5.00^{\pm 0.00}$ $5.00^{\pm 0.00}$ $4.00^{\pm 0.71}$ $4.00^{\pm 0.00}$ $4.50^{\pm 0.50}$ $4.75^{\pm 0.43}$ $6.00^{\pm 0.00}$ $6.00^{\pm 0.00}$ $6.00^{\pm 0.00}$
GPTScore(davinci-002)[25] $8.00^{\pm 0.00}$ $6.25^{\pm 0.43}$ $6.00^{\pm 0.71}$ $6.00^{\pm 0.00}$ $6.50^{\pm 0.50}$ $6.25^{\pm 0.43}$ $8.00^{\pm 0.00}$ $8.25^{\pm 0.83}$ $8.25^{\pm 1.48}$
PandaLM[52] $5.00^{\pm 0.00}$ $5.50^{\pm 0.50}$ $6.00^{\pm 0.00}$ $7.00^{\pm 0.00}$ $7.00^{\pm 0.71}$ $7.25^{\pm 0.43}$ $6.00^{\pm 0.00}$ $5.75^{\pm 0.43}$ $5.50^{\pm 0.50}$
PRD[31] $8.00^{\pm 0.00}$ $8.75^{\pm 0.43}$ $9.25^{\pm 0.83}$ $8.00^{\pm 0.00}$ $8.25^{\pm 0.43}$ $7.75^{\pm 0.83}$ $8.50^{\pm 0.00}$ $8.25^{\pm 0.83}$ $8.25^{\pm 0.43}$
PRE[17] $9.00^{\pm 0.00}$ $10.25^{\pm 0.43}$ $10.00^{\pm 0.87}$ $8.00^{\pm 0.00}$ $8.50^{\pm 0.50}$ $8.25^{\pm 0.83}$ $8.00^{\pm 0.00}$ $8.00^{\pm 0.00}$ $8.25^{\pm 0.43}$
PiCO (Ours) $\bm{10.00^{\pm 0.00}}$ $\bm{10.25^{\pm 0.71}}$ $\bm{10.50^{\pm 0.43}}$ $\bm{8.75^{\pm 0.43}}$ $\bm{8.75^{\pm 0.87}}$ $\bm{9.00^{\pm 1.22}}$ $\bm{9.00^{\pm 0.00}}$ $\bm{8.75^{\pm 0.43}}$ $\bm{8.50^{\pm 0.50}}$

### v3

@@ 3 Experiments
Datasets. To validate the effectiveness of the proposed approach, we perform experiments on Chatbot Arena (Zheng et al., 2023), MT-Bench (Zheng et al., 2023), and AlpacaEval (Li et al., 2023b). The MT-Bench dataset assesses six LLMs’ responses to 80 multi-category questions. The Chatbot Arena Conversations Dataset, with 33K conversations from 13K IPs during April-June 2023, evaluates real dialogue performance. AlpacaEval dataset integrates 805 evaluations from diverse tests (e.g., Self-Instruct (Wang et al., 2022), OASST, Anthropic’s helpful (Bai et al., 2022), Vicuna (Chiang et al., 2023) and Koala (Geng et al., 2023a) test sets) to align evaluations real-world interactions (Dubois et al., 2023). These datasets are collected by crowdsourcing platforms from human feedback, so they have a ground-truth ranking LLMs $\mathcal{R}^{*}$ to measure the alignment performance of different evaluation methods.
LLMs Pool. In our experiments, we employ 15 LLMs with diverse architectures to construct the LLMs pool, including GPT-3.5-Turbo (OpenAI, 2022), WizardLM-13B (Xu et al., 2023), Guanaco-33B (gua, 2023), Vicuna-7B (Chiang et al., 2023), Vicuna-13B (Chiang et al., 2023), Koala-13B (Geng et al., 2023b), Mpt-7B (Team, 2023), gpt4all-13B (Anand et al., 2023), ChatGLM-6B (Zeng et al., 2022), Oasst-sft-4-pythia-12B (Contributors, 2023), FastChat-T5-3B (Zheng et al., 2023), StableLM-7B (AI, 2023), Dolly-12B (Conover et al., 2023), LLaMA-13B (Touvron et al., 2023), Alpaca-13B (Taori et al., 2023). All models use the same prompt template, which can be found in Appendix C.
Table 2: Comparison of all methods on three datasets under data volumes of 1, 0.7, and 0.4, where the top value is highlighted by bold font. Higher $S$ and $\tau$ scores indicate better performance, while a lower $H$ score signifies improved performance.
@@
Spearman’s Rank Correlation Coefficient $S(\uparrow)$
Majority Voting $0.76^{\pm 0.00}$ $0.75^{\pm 0.01}$ $0.73^{\pm 0.03}$ $0.73^{\pm 0.00}$ $0.77^{\pm 0.01}$ $0.75^{\pm 0.01}$ $0.80^{\pm 0.00}$ $0.79^{\pm 0.01}$ $0.78^{\pm 0.01}$
Rating Voting $0.74^{\pm 0.00}$ $0.72^{\pm 0.02}$ $0.71^{\pm 0.02}$ $0.80^{\pm 0.00}$ $0.78^{\pm 0.02}$ $0.74^{\pm 0.03}$ $0.77^{\pm 0.00}$ $0.77^{\pm 0.01}$ $0.78^{\pm 0.01}$
GPTScore(flan-t5-xxl) $-0.09^{\pm 0.00}$ $-0.09^{\pm 0.01}$ $-0.12^{\pm 0.02}$ $0.05^{\pm 0.00}$ $0.01^{\pm 0.07}$ $0.04^{\pm 0.09}$ $0.34^{\pm 0.00}$ $0.34^{\pm 0.00}$ $0.34^{\pm 0.01}$
GPTScore(davinci-002) $0.15^{\pm 0.00}$ $0.13^{\pm 0.02}$ $-0.02^{\pm 0.14}$ $0.52^{\pm 0.00}$ $0.42^{\pm 0.05}$ $0.45^{\pm 0.05}$ $0.76^{\pm 0.00}$ $0.77^{\pm 0.07}$ $0.75^{\pm 0.06}$
PandaLM $0.43^{\pm 0.00}$ $0.44^{\pm 0.03}$ $0.44^{\pm 0.10}$ $0.50^{\pm 0.00}$ $0.50^{\pm 0.08}$ $0.52^{\pm 0.17}$ $0.57^{\pm 0.00}$ $0.55^{\pm 0.01}$ $0.48^{\pm 0.08}$
PRD $0.84^{\pm 0.00}$ $0.84^{\pm 0.00}$ $0.82^{\pm 0.03}$ $0.86^{\pm 0.00}$ $0.84^{\pm 0.03}$ $0.81^{\pm 0.03}$ $0.81^{\pm 0.00}$ $0.81^{\pm 0.01}$ $0.81^{\pm 0.02}$
PRE $0.86^{\pm 0.00}$ $0.86^{\pm 0.01}$ $0.86^{\pm 0.01}$ $0.86^{\pm 0.00}$ $0.84^{\pm 0.03}$ $0.82^{\pm 0.04}$ $0.83^{\pm 0.00}$ $0.81^{\pm 0.01}$ $0.83^{\pm 0.02}$
Claude-3 (API) $0.90^{\pm 0.01}$ $0.88^{\pm 0.03}$ $0.87^{\pm 0.04}$ $0.85^{\pm 0.06}$ $0.82^{\pm 0.08}$ $0.80^{\pm 0.07}$ $0.79^{\pm 0.03}$ $0.78^{\pm 0.02}$ $0.75^{\pm 0.04}$
PiCO (Ours) $\bm{0.90^{\pm 0.00}}$ $\bm{0.89^{\pm 0.01}}$ $\bm{0.89^{\pm 0.01}}$ $\bm{0.89^{\pm 0.01}}$ $\bm{0.89^{\pm 0.01}}$ $\bm{0.84^{\pm 0.11}}$ $\bm{0.84^{\pm 0.00}}$ $\bm{0.83^{\pm 0.03}}$ $\bm{0.85^{\pm 0.01}}$
Kendall’s Rank Correlation Coefficient $\tau(\uparrow)$
Majority Voting $0.58^{\pm 0.00}$ $0.56^{\pm 0.02}$ $0.52^{\pm 0.05}$ $0.56^{\pm 0.00}$ $0.61^{\pm 0.02}$ $0.60^{\pm 0.02}$ $0.62^{\pm 0.00}$ $0.60^{\pm 0.02}$ $0.58^{\pm 0.02}$
Rating Voting $0.54^{\pm 0.00}$ $0.53^{\pm 0.02}$ $0.52^{\pm 0.02}$ $0.58^{\pm 0.00}$ $0.57^{\pm 0.02}$ $0.54^{\pm 0.01}$ $0.58^{\pm 0.00}$ $0.57^{\pm 0.01}$ $0.57^{\pm 0.02}$
GPTScore(flan-t5-xxl) $-0.06^{\pm 0.00}$ $-0.06^{\pm 0.02}$ $-0.09^{\pm 0.02}$ $-0.05^{\pm 0.00}$ $-0.07^{\pm 0.05}$ $-0.02^{\pm 0.06}$ $0.25^{\pm 0.00}$ $0.26^{\pm 0.01}$ $0.26^{\pm 0.01}$
GPTScore(davinci-002) $0.20^{\pm 0.00}$ $0.23^{\pm 0.02}$ $0.03^{\pm 0.11}$ $0.36^{\pm 0.00}$ $0.30^{\pm 0.05}$ $0.31^{\pm 0.05}$ $0.60^{\pm 0.08}$ $0.61^{\pm 0.05}$ $0.59^{\pm 0.08}$
PandaLM $0.30^{\pm 0.00}$ $0.31^{\pm 0.03}$ $0.31^{\pm 0.07}$ $0.39^{\pm 0.00}$ $0.37^{\pm 0.06}$ $0.40^{\pm 0.12}$ $0.41^{\pm 0.00}$ $0.39^{\pm 0.02}$ $0.32^{\pm 0.05}$
PRD $0.68^{\pm 0.00}$ $0.69^{\pm 0.01}$ $0.67^{\pm 0.03}$ $0.68^{\pm 0.06}$ $0.66^{\pm 0.02}$ $0.63^{\pm 0.03}$ $0.64^{\pm 0.00}$ $0.63^{\pm 0.03}$ $0.63^{\pm 0.02}$
PRE $0.71^{\pm 0.00}$ $0.73^{\pm 0.02}$ $0.72^{\pm 0.02}$ $0.68^{\pm 0.00}$ $0.68^{\pm 0.02}$ $0.65^{\pm 0.03}$ $0.64^{\pm 0.00}$ $0.66^{\pm 0.01}$ $0.66^{\pm 0.03}$
Claude-3 (API) $0.76^{\pm 0.04}$ $0.72^{\pm 0.05}$ $0.70^{\pm 0.07}$ $0.67^{\pm 0.07}$ $0.66^{\pm 0.11}$ $0.61^{\pm 0.10}$ $0.64^{\pm 0.06}$ $0.61^{\pm 0.04}$ $0.66^{\pm 0.06}$
PiCO (Ours) $0.77^{\pm 0.00}$ $0.76^{\pm 0.01}$ $0.77^{\pm 0.02}$ $0.72^{\pm 0.01}$ $0.72^{\pm 0.03}$ $0.70^{\pm 0.12}$ $0.68^{\pm 0.00}$ $0.66^{\pm 0.04}$ $0.67^{\pm 0.02}$
Permutation Entropy $H(\downarrow)$
Majority Voting $1.27^{\pm 0.05}$ $1.30^{\pm 0.03}$ $1.36^{\pm 0.06}$ $1.37^{\pm 0.03}$ $1.30^{\pm 0.06}$ $1.27^{\pm 0.04}$ $1.26^{\pm 0.02}$ $1.28^{\pm 0.03}$ $1.29^{\pm 0.03}$
Rating Voting $1.39^{\pm 0.02}$ $1.43^{\pm 0.03}$ $1.42^{\pm 0.07}$ $1.32^{\pm 0.03}$ $1.35^{\pm 0.04}$ $1.38^{\pm 0.04}$ $1.34^{\pm 0.03}$ $1.37^{\pm 0.03}$ $1.34^{\pm 0.08}$
GPTScore(flan-t5-xxl) $1.68^{\pm 0.01}$ $1.68^{\pm 0.02}$ $1.65^{\pm 0.02}$ $1.72^{\pm 0.02}$ $1.70^{\pm 0.02}$ $1.68^{\pm 0.03}$ $1.55^{\pm 0.02}$ $1.57^{\pm 0.03}$ $1.60^{\pm 0.01}$
GPTScore(davinci-002) $1.54^{\pm 0.02}$ $1.64^{\pm 0.02}$ $1.68^{\pm 0.05}$ $1.51^{\pm 0.02}$ $1.61^{\pm 0.01}$ $1.61^{\pm 0.04}$ $1.25^{\pm 0.02}$ $1.23^{\pm 0.08}$ $1.26^{\pm 0.14}$
PandaLM $1.65^{\pm 0.01}$ $1.64^{\pm 0.02}$ $1.63^{\pm 0.05}$ $1.55^{\pm 0.03}$ $1.59^{\pm 0.05}$ $1.52^{\pm 0.08}$ $1.56^{\pm 0.01}$ $1.58^{\pm 0.01}$ $1.64^{\pm 0.05}$
PRD $1.15^{\pm 0.04}$ $1.12^{\pm 0.05}$ $1.13^{\pm 0.06}$ $1.15^{\pm 0.05}$ $1.17^{\pm 0.06}$ $1.23^{\pm 0.04}$ $1.21^{\pm 0.04}$ $1.22^{\pm 0.06}$ $1.23^{\pm 0.07}$
PRE $1.07^{\pm 0.01}$ $1.03^{\pm 0.03}$ $1.06^{\pm 0.04}$ $1.17^{\pm 0.04}$ $1.13^{\pm 0.05}$ $1.19^{\pm 0.05}$ $1.18^{\pm 0.03}$ $1.21^{\pm 0.04}$ $1.15^{\pm 0.05}$
@@
  • The wisdom of the crowds: The two methods that perform LLMs evaluation based on the wisdom of the crowds (Surowiecki, 2005; Budescu & Chen, 2015; Weller, 2007) are compared in this experiment. 1) Majority Voting (Surowiecki, 2005): Multiple review models vote for the better answer for the same response pair, and the model with the most votes gets 1 score; 2) Rating Voting (Allahbakhsh & Ignjatovic, 2012): Multiple review models also vote on the same response pair, and the number of votes obtained is the score.
  • State-of-the-art methods: The four recent SOTA methods of using either single or multiple models for self-evaluation are compared in this experiment. PandaLM (Wang et al., 2023): It is a fine-tuned language model based on Llama-7b designed for the preference judgment tasks to evaluate and optimize LLMs. GPTScore (Fu et al., 2023): It employs generative pre-trained models to assess the quality of generated text. It calculates the likelihood that the text was generated in response to specific instructions and context, indicative of high quality. In our implementation, GPT-3 (davinci-002) and flan-t5-xxl serve as the base models. PRD (Li et al., 2023a): It transforms the LLMs win rates into weights for competitive ranking, while evaluating each LLM based on its preference for all possible pairs of answers, enabling a tournament-style ranking system. PRE (Chu et al., 2024): It employs a supervised process to evaluate LLMs using a qualification exam, aggregates their scores based on accuracy, and assigns weights accordingly. Claude-3 (API): Another SOTA closed-source LLM developed by Anthropic. PiCO (Ours): the proposed approach in this paper.
Metrics. For all experiments, we employ three popular rank-based metrics to evaluate the aforementioned experimental setups and our PiCO method: Spearman’s Rank Correlation Coefficient $S(\uparrow)$ (Lehman et al., 2013), Kendall’s Rank Correlation Coefficient $\tau(\uparrow)$ (Kendall, 1938) and Permutation Entropy $H(\downarrow)$ (Bandt & Pompe, 2002). The details of these metrics can be found in the Appendix A. Moreover, we perform the experiments for 4 runs and record the average results over 4 seeds ($seed=1,2,3,4$).